# State Space Models for Event Cameras

### v2

@@ 4.3 Ablation study
In this section we investigate different SSMs and their initializations with and without bandlimiting parameter for event-based vision to address the problem of inference at different frequencies (4.3.1). After that, we study the importance of SSM layers in various stages during training of the model (4.3.2). Finally, we evaluate models on differing frequencies with two proposed strategies (4.3.3).
@@ 4.3.1 SSMs: initializations & bandlimiting
This section presents an ablation study focusing on the performance impact of various SSM variants and their initialization strategies in conjunction with the bandlimiting parameter $\alpha$. The SSM variants under consideration are S4 [19], S4D [18], and the more recent S5 [41]. The analysis is conducted on the Gen1 [8] validation dataset, with the mean Average Precision (mAP) serving as the performance metric. Table 3 provides a comprehensive view of how different initialization strategies (legS, inv, lin) introduced in [18] and values of $\alpha\in\{0,0.5,1\}$ influence model performance. The S4-legS model, not equipped for bandlimiting in non-diagonal matrix scenarios, achieves a baseline mAP of 46.66 at $\alpha=0$. The S4D variants demonstrate a diverse performance spectrum. The S4D-legS variant, particularly at $\alpha=0.5$, achieves the highest mAP of 47.33 among the S4D models, also maintaining the best average mAP of 46.92. The S4D-inv and S4D-lin variants show less favorable outcomes, with mAPs peaking at 46.23 and 46.02, respectively, for $\alpha=0.5$.
Notably, the S5 model variants exhibit an improvement over their predecessors. The S5-legS variant achieves the highest mAP of 48.48 at $\alpha=0.5$, and also records the best average mAP of 48.27. This result is not only the best in its category but also the best overall. The S5-inv and S5-lin models also demonstrate comparable performance, with the former reaching its peak mAP of 47.43 at $\alpha=0.5$.
@@ 4.3.2 SSM Utilization Analysis
Table 4 shows the outcomes of these manipulations on the Gen1 validation dataset. $S4D_{.5}$ represents S4D model, and $S5_{.5}$ is S5 model with $\alpha=0.5$. The data clearly indicate that the complete removal of SSMs from the network leads to a highest decrease in detection performance. This underscores the pivotal role of SSMs in enhancing the model’s capability. On the other hand, initiating the use of SSMs from the fourth stage onwards consistently enhances performance, suggesting a critical threshold for the impact of temporal information processing in the later stages of the network. Another intriguing observation is the performance boost obtained by incorporating an SSM at the very initial stage. This suggests that the early integration of temporal information is also beneficial for the overall detection performance. As a result of these findings, our preferred configuration includes the SSM component right from the initial stage, thereby leveraging the advantages of temporal information processing throughout the entire network. Noteworthy is the observation that the performance drop when employing SSMs is less pronounced than with RVT, suggesting our approach’s superior robustness to temporal aggregation from certain stages compared to RVT.
@@ 4.3.3 Evaluation at different frequencies
In this section, we delve into the comparative performance analysis of three models: RVT [13], S4D [18], and S5 [41], across various frequency ranges. This evaluation is crucial in determining the robustness and adaptability of these models under diverse inference frequencies, specifically at 20 Hz, 40 Hz, 80 Hz, 100 Hz, and 200 Hz. Our model is trained with event representations formed based on 50 ms time windows that correspond to 20 Hz sampling frequency. Evaluation is done at this frequency, and others above mentioned. With RVT, there are no changes when evaluating at different frequencies, while with SSMs rate $r$ is halved for the double the inference frequency etc.
@@ 20 Hz 40 Hz 80 Hz 100 Hz 200 Hz
[…]
The analysis reveals that both the $H_{2}$ norm and bandlimiting strategies with $\alpha=0.5$ offer comparable performance across the assessed frequency ranges. However, a slight edge is observed with the $H_{2}$ norm, particularly at very high frequencies. This marginal superiority can be attributed to the fact that $H_{2}$ norm approach does not explicitly mask output’s ($C$) matrix columns.
@@ 13 SSM-only and SSM-ConvNext models
To study the performance of SSM-only models, we replaced the attention block with a 2D-SSM block and trained it on Gen1 and 1 Mpx datasets. We obtain 46.1 mAP and 45.74 mAP which is lower than the original S5-ViT-B model’s performance, as it can be seen in Table 6. To evaluate a CNN-based backbone, we replace the attention block with the ConvNext block [31] showing that ViT achieves better performance than the CNN structure.
[…]
S5-ViT-B 47.40 47.20

### v3

@@ 4.3 Ablation study
In this section, we investigate different SSMs and their initializations with and without bandlimiting parameter for event-based vision to address the problem of inference at different frequencies (4.3.1). After that, we study the importance of SSM layers in various stages during the training of the model (4.3.2). Finally, we evaluate models on differing frequencies with two proposed strategies (4.3.3).
@@ 4.3.1 SSMs: initializations & bandlimiting
This section presents an ablation study focusing on the performance impact of various SSM variants and their initializations in conjunction with the bandlimiting parameter $\alpha$. The SSM variants under consideration are S4 [19], S4D [18], and the more recent S5 [41]. The analysis is conducted on the Gen1 [8] validation dataset, with the mAP serving as the performance metric. Table 3 provides a comprehensive view of how different initialization strategies (legS, inv, lin) introduced in [18] and values of $\alpha\in\{0,0.5,1\}$ influence model performance. The S4-legS model, not equipped for bandlimiting in non-diagonal matrix scenarios, achieves a baseline mAP of 46.66 at $\alpha=0$. The S4D variants demonstrate a diverse performance spectrum. The S4D-legS variant, particularly at $\alpha=0.5$, achieves the highest mAP of 47.33 among the S4D models, also maintaining the best average mAP of 46.92. The S4D-inv and S4D-lin variants show less favorable outcomes, with mAPs peaking at 46.23 and 46.02, respectively, for $\alpha=0.5$.
Notably, the S5 model variants exhibit an improvement over their predecessors. The S5-legS variant achieves the highest mAP of 48.48 at $\alpha=0.5$ and also records the best average mAP of 48.27. This result is not only the best in its category but also the best overall. The S5-inv and S5-lin models also demonstrate comparable performance, with the former reaching its peak mAP of 47.43 at $\alpha=0.5$.
@@ 4.3.2 SSM Utilization Analysis
Table 4 shows the outcomes of these manipulations on the Gen1 validation dataset. $S4D_{.5}$ represents S4D model, and $S5_{.5}$ is S5 model with $\alpha=0.5$. The data clearly indicate that the complete removal of SSMs from the network leads to the highest decrease in detection performance. This underscores the critical role of SSMs in enhancing the model’s capability. On the other hand, initiating the use of SSMs from the fourth stage onwards consistently enhances performance, suggesting a critical threshold for the impact of temporal information processing in the later stages of the network. Another intriguing observation is the performance boost obtained by incorporating an SSM at the very initial stage. This suggests that the early integration of temporal information is also beneficial for the overall detection performance. As a result of these findings, our preferred configuration includes the SSM component right from the initial stage, thereby leveraging the advantages of temporal information processing throughout the entire network. Noteworthy is the observation that the performance drop when employing SSMs is less pronounced than with RVT, suggesting our approach’s superior robustness to temporal aggregation from certain stages compared to RVT.
@@ 4.3.3 Evaluation at different frequencies
In this section, we compare the performance of three models: RVT [13], S4D [18], and S5 [41], across various frequency ranges. This evaluation is crucial in determining the robustness and adaptability of these models under diverse inference frequencies, specifically at 20 Hz, 40 Hz, 80 Hz, 100 Hz, and 200 Hz. Our model is trained with event representations formed based on 50 ms time windows that correspond to a 20 Hz sampling frequency. Evaluation is done at this frequency, and others above mentioned. With RVT, there are no changes when evaluating at different frequencies, while with SSMs rate $r$ is halved for double the inference frequency, etc.
@@ 20 Hz 40 Hz 80 Hz 100 Hz 200 Hz
Each model in Table 5 is assessed across datasets and model sizes, including Gen1Base, Gen1Small, and 1 MpxBase. $Base$ and $Small$ represent two variants of models on the Gen1 and 1 Mpx datasets, the base one being the larger one with 16.5M parameters for the S4D model and 18.2M parameters for the S5 model (as presented in Table 1 parameters’ column), and small one with 8.8M parameters for S4D and 9.7M parameters for S5 model. The focal point of this analysis is the performance drop, calculated as the average difference between the original performance at 20 Hz and performances at higher frequencies. A notable aspect of this study is the inherent advantage of the S4D and S5 models due to their incorporation of a learnable timescale parameter. This significantly enhances their adaptability, allowing them to dynamically adjust to varying frequencies. They are further analyzed based on the different operational strategies: the $H_{2}$ norm and bandlimiting with $\alpha=0.5$. The inclusion of the learnable timescale parameter in these models underscores their capability to maintain performance across a wide range of frequencies.
The analysis reveals that both the $H_{2}$ norm and bandlimiting strategies with $\alpha=0.5$ offer comparable performance across the assessed frequency ranges. However, a slight edge is observed with the $H_{2}$ norm, particularly at very high frequencies. This marginal superiority can be attributed to the fact that the $H_{2}$ norm approach does not explicitly mask the output’s ($C$) matrix columns.
@@ 13 SSM-only and SSM-ConvNext models
To study the performance of SSM-only models, we replaced the attention block with a 2D-SSM block and trained it on Gen1 and 1 Mpx datasets. We obtain 46.10 mAP and 45.74 mAP which is lower than the original S5-ViT-B model’s performance, as it can be seen in Table 6. To evaluate a CNN-based backbone, we replace the attention block with the ConvNext block [31] showing that ViT achieves better performance than the CNN structure.
Model mAPGen1 mAP1Mpx
S5-ViT-B 47.71 47.80